Prove that Hilbert space is seperable
1
Expert's answer
2020-05-12T18:28:55-0400

Theorem: A Hilbert space H is separable (that is, has a countable dense subset) if and only if it has one countable orthonormal basis if and only if every orthonormal basis for H is countable. 


Proof: "\implies": Suppose H is a separable and consider any orthonormal basis "S\subset H". Then if "\phi_1,\phi_2\in S" and "\phi_1{=}\mathllap{/\,}\phi_2" we conclude "||\phi_1-\phi_2||=\sqrt{(\phi_1-\phi_2,\phi_1-\phi_2)}=\sqrt{2}". Thus "S\subset H" has the discrete topology.

Now "H" a separable metric space "\implies S" is a separable metric space "\implies S" has a countable dense subset. But "S" is the only subset of itself that is dense in "S" (since "S" has the discrete topology) and so "S" must be countable.


"\impliedby": Suppose for the converse that there is one countable orthonormal basis "S=\{\phi_1,\phi_2,...\}" for "H". We look at the case "\Kappa=\R" ("H" is a real Hilbert space) first. It is quite easy to check that the sets

"D_n=\{\sum""n\atop j=1" "q_j \phi_j:q_j\in Q" for "1\leq j\leq n\}"

are each countable and that their union "D=\bigcup""\infin\atop n=1""D_n" is countable and dense in "H". The closure of each "D_n" is easily seen to be the "\R"-linear span of "\phi_1,\phi_2,...,\phi_n" and so the closure of "D" includes all finite linear combinations "\sum""n\atop j=1""x_j\phi_j". But, each "x\in H" is a limit of such finite linear combinations. Hence the closure of "D" is all of "H". As "D" is countable, this shows that "H" must be separable.

In the complex case ("H" is a Hilbert space over "K=\Complex" we must take "q_j\in Q+iQ" instead, so that we can get all finite "\Complex"-linear combinations of the "\phi_j" n the closure of "D" (and there is no other difference in the proof). "\blacksquare"